# Accuracy of the Ensemble Kalman Filter in the Near-Linear Setting

### v1

@@ Abstract
The filtering distribution captures the statistics of the state of a dynamical system from partial and noisy observations. Classical particle filters provably approximate this distribution in quite general settings; however they behave poorly for high dimensional problems, suffering weight collapse. This issue is circumvented by the ensemble Kalman filter which is an equal-weight interacting particle system. However, this finite particle system is only proven to approximate the true filter in the linear Gaussian case. In practice, however, it is applied in much broader settings; as a result, establishing its approximation properties more generally is important. There has been recent progress in the theoretical analysis of the algorithm, establishing stability and error estimates in non-Gaussian settings, but the assumptions on the dynamics and observation models rule out the unbounded vector fields that arise in practice and the analysis applies only to the mean field limit of the ensemble Kalman filter. The present work establishes error bounds between the filtering distribution and the finite particle ensemble Kalman filter when the model exhibits linear growth.
@@ 1.1 Literature Review, Contributions and Outline
Algorithms for filtering employ noisy observations arising from a possibly random dynamical system to estimate the distribution of the state of the dynamical system conditional on the observations. The Kalman filter [kalman1960new] determines the filtering distribution exactly for linear Gaussian dynamics and observations. The extended Kalman filter was developed as an extension of the Kalman filtering technique to nonlinear problems and is based on a linearization approximation; see [jazwinski2007stochastic] and [anderson2012optimal]. The linearization approximation leads to an inexact distribution for nonlinear problems, and furthermore requires evaluation of covariance matrices, making the methodology impractical for large-scale geophysical applications [ghil]. Particle filters or sequential Monte Carlo methods [doucet2001sequential, chopin:20] offer an alternative methodology for nonlinear filtering problems, allowing recovery of the exact filtering distribution in the large particle limit. However, this class of methods scales poorly with dimension, and in particular suffer weight collapse, making their application to geophysical problems prohibitive; see for example [snyder2008obstacles, bickel2008sharp, rebeschini2015can, agapiou2017importance].
The ensemble Kalman filter was introduced in the seminal paper [evensen1994sequential]. Its success stems from the low-rank approximation of large covariances, cheaply computed using an ensemble of particles and allowing it to be deployed in geophysical applications. However analysis of its accuracy, in relation to the true filtering distribution, remains in its infancy. The papers [le2009large, mandel2011convergence] studied this issue in the linear Gaussian setting where the mean field limit of the Kalman filter is exact; they demonstrate that the ensemble Kalman filter may be viewed as an interacting particle system approximation of the mean field limit, and establish Monte Carlo type error bounds. The recent article [calvello2025ensemble] overviews the formulation of ensemble Kalman methods using mean field dynamical systems and provides a platform from which the analyses of [le2009large, mandel2011convergence] may be generalized beyond the linear Gaussian setting. In the recent paper [carrillo2022ensemble] the authors establish stability properties of the mean field ensemble Kalman filter and use them to prove accuracy of the filter in a near-Gaussian setting. However the paper does not consider finite particle approximations of the mean field, and the conditions on the dynamics-observation model require boundedness of the vector fields arising. In this paper we address both these issues, establishing error bounds between the finite particle ensemble Kalman filter and the true filtering distribution in settings where the dynamics and observation vector fields grow linearly. We make the following contributions.
@@
In going beyond the work in [carrillo2022ensemble], the current paper simultaneously addresses a more applicable problem class, by allowing linear growth in the dynamics and observations operators, and confronts the substantial technical challenges which arise from doing so. Bounds on moments of the filtering distribution and mean field ensemble Kalman filter must be established; these bounds grow in the number of iterations which is in contrast to [carrillo2022ensemble] where the $L^{\infty}$ bounds on model and observation operator ensure a uniform bound on moments. A further challenge is presented by the need to establish stability bounds for the conditioning map, giving rise to the true filter, and the transport map giving rise to the ensemble Kalman filter. These results exhibit dependence on moments in the stability constants and require control of the growth given by the dynamics; this is again in contrast to [carrillo2022ensemble] where the $L^{\infty}$ bounds allow for a uniform control.
@@ 1.2 Notation
We let $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ denote the $C^{0,1}$ semi-norm, namely the Lipschitz constant.
@@
The notations $\mathcal{M}(\mu)$ and $\mathcal{C}(\mu)$ denote respectively the mean and covariance under $\mu$:
@@
The notation $\mathcal{P}_{R}(^{n})$ for $R\geqslant 1$ refers to the subset of $\mathcal{P}(^{n})$ of probability measures whose mean and covariance satisfy the bound
$\lvert\mathcal{M}(\mu)\rvert\leqslant R,\qquad\frac{1}{R^{2}}I_{n}\preccurlyeq\mathcal{C}(\mu)\preccurlyeq R^{2}I_{n}.$ (1.2)
@@
For a probability measure $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ associated with random variable $(u,y)\in^{d_{u}}\times^{d_{y}}$ we use the notation $\mathcal{M}^{u}(\pi)$, $\mathcal{M}^{y}(\pi)$ for the means of the marginal distributions, and the notation $\mathcal{C}^{uu}(\pi)$, $\mathcal{C}^{uy}(\pi)$ and $\mathcal{C}^{yy}(\pi)$ for the blocks of the covariance matrix $\mathcal{C}(\pi)$. That is to say,
@@
For $h\colon^{d_{u}}\to^{d_{y}}$ we also define ${\mathcal{C}}^{hh}(\pi)$ to be the covariance of vector $h(u)$, for $u$ distributed according to the marginal of $\pi$ on $u$, and ${\mathcal{C}}^{uh}(\pi)$ to be the covariance between $u$ and $h(u)$.
We also introduce the Gaussian projection operator $\mathsf{G}\colon\mathcal{P}(^{n})\to\mathcal{G}(^{n})$ given by $\mathsf{G}\mu=\mathcal{N}\bigl{(}\mathcal{M}(\mu),\mathcal{C}(\mu)\bigr{)}.$ We refer to $\mathsf{G}$ as a projection because $\mathsf{G}\mu$ is the Gaussian distribution closest to $\mu$ with respect to $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ [MR2247587], where $\mathrm{KL}(\mu\|\nu)$ is the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$. Note that $\mathsf{G}$ defines a nonlinear mapping. Throughout this work, we employ the following weighted total variation distance, shown to be useful in the analysis of ensemble Kalman methods in [carrillo2022ensemble]:
@@
Let $g\colon^{n}\to[1,\infty)$ denote $g(v):=1+\lvert v\rvert^{2}$. We define the weighted total variation metric $d_{g}\colon\mathcal{P}(^{n})\times\mathcal{P}(^{n})\to$
$d_{g}(\mu_{1},\mu_{2})=\sup_{\lvert f\rvert\leqslant g}\bigl{\lvert}\mu_{1}[f]-\mu_{2}[f]\bigr{\rvert},$
where the supremum is over all functions $f\colon^{n}\to$ which are bounded from above by $g$ pointwise and in absolute value.
@@
  • Unlike the usual total variation distance, the weighted metric in Definition 1.1 enables control of the differences $\lvert\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\rvert$ and $\lVert\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\rVert$. This is the content of Lemma B.6, proved in the appendix, which is used in the proof of a key auxiliary result (Lemma B.2).
@@ 1.3 Filtering Distribution
We consider the following stochastic dynamics and data model
@@
Here $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}$ is the unknown state, evolving in ${}^{d_{u}}$, and $\{y_{j}\}_{j\in\llbracket 1,J\rrbracket}$ are the observations, evolving in ${}^{d_{y}}$. We assume that the initial state is distributed according to a Gaussian distribution $u_{0}\sim\mathcal{N}(m_{0},C_{0})$ and that the following independence assumption is satisfied:
@@
The filtering distribution $\mu_{j}$ is the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}:=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ of the data process up to step $j$. Data $Y_{j}^{\dagger}$ may be thought of as arising from a realization of (1.4); but the case of model misspecification, where $Y_{j}^{\dagger}$ does not necessarily arise from (1.4), is also of interest.
It is well-known [MR3363508, reich2015probabilistic] that the true filtering distribution evolves according to
$\mu_{j+1}=\mathsf{L}_{j}\mathsf{P}\mu_{j}.$ (1.6)
where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps on probability measures, referred to respectively as the prediction and analysis steps in the data assimilation community [asch2016data]. The prediction operator $\mathsf{P}$ is linear and defined by the Markov kernel associated with the stochastic dynamics (1.4a). Its action on a probability measure with Lebesgue density $\mu$ reads
$\mathsf{P}\mu(u)=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\mu(v)\,\mathrm{d}v.$ (1.7)
The operator $\mathsf{L}_{j}$ is a nonlinear map which formalizes the incorporation of the new datum $y^{\dagger}_{j+1}$ using Bayes’ theorem. Its action reads
$\mathsf{L}_{j}\mu(u)=\frac{\displaystyle\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u)}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(U)\,\mathrm{d}U}$ (1.8)
The operator $\mathsf{L}_{j}$ effects a reweighting of the measure to which it applies, with more weight assigned to the state values that are consistent with the observation. It is convenient in this work to decompose the analysis map $\mathsf{L}_{j}$ into the composition $\mathsf{B}_{j}\mathsf{Q}$, where the operators $\mathsf{Q}\colon\mathcal{P}(^{d_{u}})\to\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and $\mathsf{B}_{j}\colon\mathcal{P}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ are defined by
$\mathsf{Q}\mu(u,y)=\frac{1}{\sqrt{(2\pi)^{d_{y}}\det\Gamma}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u).$ (1.9)
and
$\mathsf{B}_{j}\mu(u)=\frac{\mu(u,y_{j+1}^{\dagger})}{\displaystyle\int_{{}^{d_{u}}}\mu(U,y_{j+1}^{\dagger})\,\mathrm{d}U}.$ (1.10)
The operator $\mathsf{Q}$ maps a probability measure with density $\mu$ into the density associated with the joint random variable $\bigl{(}U,h(U)+\eta\bigr{)}$, where $U\sim\mu$ is independent of $\eta\sim\mathcal{N}(0,\Gamma)$. The operator $\mathsf{B}_{j}$ performs conditioning on the data $y^{\dagger}_{j+1}$. Map $\mathsf{Q}$ is linear whilst $\mathsf{B}_{j}$ is nonlinear. We may thus write (1.6) in the form
@@ 2 The Ensemble Kalman Filter
In Subsection 2.1 we define the specific version of the mean field ensemble Kalman filter that we analyze here; other versions may be found in [calvello2025ensemble] and will be amenable to similar analyses. In Subsection 2.2 we prove our main stability theorem, showing that the error between the true filter and its Gaussian projection, on the joint space of state and observations, may be used to control the error between the true filter and the mean field ensemble Kalman filter. In Subsection 2.3 we prove a corollary to this theorem, establishing that the mean field ensemble Kalman filter accurately approximates the true filter for a specific class of non-Gaussian problems. In Subsection 2.4 we deploy the results of the two preceding subsections to state a theorem quantifying the error between the finite particle ensemble Kalman filter itself and the true filtering distribution. All of the theorems allow for linear growth of the vector fields defining the dynamics and the observation processes.

### v2

@@ Abstract
The filtering distribution captures the statistics of the state of a dynamical system from partial and noisy observations. Classical particle filters provably approximate this distribution in quite general settings; however they behave poorly for high dimensional problems, suffering weight collapse. This issue is circumvented by the ensemble Kalman filter which is an equal-weight interacting particle system. However, this finite particle system is only proven to approximate the true filter in the linear Gaussian case. In practice, however, it is applied in much broader settings; as a result, establishing its approximation properties more generally is important. There has been recent progress in the theoretical analysis of the algorithm, establishing stability and error estimates in non-Gaussian settings, but the assumptions on the dynamics and observation models rule out the unbounded vector fields that arise in practice and the analysis applies only to the mean field limit of the ensemble Kalman filter. The present work establishes error bounds between the filtering distribution and the finite particle ensemble Kalman filter when the dynamics and observation vector fields may be unbounded, allowing linear growth.
@@ 1.1 Literature Review, Contributions and Outline
Algorithms for filtering employ noisy observations arising from a possibly random dynamical system to estimate the distribution of the state of the dynamical system conditional on the observations. The Kalman filter [kalman1960new] determines the filtering distribution exactly for linear Gaussian dynamics and observations. The extended Kalman filter was developed as an extension of the Kalman filtering technique to nonlinear problems and is based on a linearization approximation; see [jazwinski2007stochastic] and [anderson2012optimal]. The linearization approximation leads to an inexact distribution for nonlinear problems, and furthermore requires evaluation of covariance matrices, making the methodology impractical for large-scale geophysical applications [ghil]. Particle filters or sequential Monte Carlo methods [doucet2001sequential, chopin:20] offer an alternative methodology for nonlinear filtering problems, allowing recovery of the exact filtering distribution in the large particle limit. However, this class of methods scales poorly with dimension, and in particular suffers weight collapse, making its application to geophysical problems prohibitive; see for example [snyder2008obstacles, bickel2008sharp, rebeschini2015can, agapiou2017importance].
The ensemble Kalman filter was introduced in the seminal paper [evensen1994sequential]. Its success stems from the low-rank approximation of large covariances, cheaply computed using an ensemble of particles, which allows the methodology to be deployed in geophysical applications. However analysis of its accuracy, in relation to the true filtering distribution, remains in its infancy. The papers [le2009large, mandel2011convergence] studied this issue in the linear Gaussian setting where the mean field limit of the Kalman filter is exact; they demonstrate that the ensemble Kalman filter may be viewed as an interacting particle system approximation of the mean field limit, and establish Monte Carlo type error bounds. The recent article [calvello2025ensemble] overviews the formulation of ensemble Kalman methods using mean field dynamical systems and provides a platform from which the analyses of [le2009large, mandel2011convergence] may be generalized beyond the linear Gaussian setting. In the recent paper [carrillo2022ensemble] the authors establish stability properties of the mean field ensemble Kalman filter and use them to prove accuracy of the filter in a near-Gaussian setting. However, the paper does not consider finite particle approximations of the mean field, and the conditions on the dynamics-observation model require boundedness of the vector fields arising. In this paper we address both these issues, establishing error bounds between the finite particle ensemble Kalman filter and the true filtering distribution in settings where the dynamics and observation vector fields grow linearly. We make the following contributions.
@@
In going beyond the work in [carrillo2022ensemble], the current paper simultaneously addresses a more applicable problem class, by allowing linear growth in the dynamics and observation operators, and confronts the substantial technical challenges which arise from doing so. Bounds on moments of the filtering distribution and mean field ensemble Kalman filter must be established; these bounds grow in the number of iterations which is in contrast to [carrillo2022ensemble] where the $L^{\infty}$ bounds on model and observation operator ensure a uniform bound on moments. A further challenge is presented by the need to establish stability bounds for the conditioning map, giving rise to the true filter, and the transport map giving rise to the ensemble Kalman filter. These results exhibit dependence on moments in the stability constants and require control of the growth given by the dynamics; this is again in contrast to [carrillo2022ensemble] where the $L^{\infty}$ bounds allow for a uniform control.
@@ 1.2 Notation
We write $|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}|_{C^{0,1}}$ for the $C^{0,1}$ semi-norm, referring in particular to the Lipschitz constant.
@@
We denote by $\mathcal{M}(\mu)$ and $\mathcal{C}(\mu)$ the mean and covariance under $\mu$, respectively:
@@
We use $\mathcal{P}_{R}(^{n})$ with $R\geqslant 1$ to denote the subset of $\mathcal{P}(^{n})$ of probability measures with mean and covariance satisfying the bounds
$\lvert\mathcal{M}(\mu)\rvert\leqslant R,\qquad\frac{1}{R^{2}}I_{n}\preccurlyeq\mathcal{C}(\mu)\preccurlyeq R^{2}I_{n},$ (1.2)
@@
For $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$ defined on the joint space of state and data associated with random variable $(u,y)\in^{d_{u}}\times^{d_{y}}$ we denote by $\mathcal{M}^{u}(\pi)$, $\mathcal{M}^{y}(\pi)$ the means of the marginal distributions, and by $\mathcal{C}^{uu}(\pi)$, $\mathcal{C}^{uy}(\pi)$ and $\mathcal{C}^{yy}(\pi)$ the blocks of the covariance matrix $\mathcal{C}(\pi)$. In particular, we write
@@
For $h\colon^{d_{u}}\to^{d_{y}}$ we also define ${\mathcal{C}}^{hh}(\pi)$ to be the covariance of vector $h(u)$, for $u$ distributed according to the marginal of $\pi$ on $u$, and ${\mathcal{C}}^{uh}(\pi)$ to be the covariance between $u$ and $h(u)$. Given this notation, we also introduce the set $\mathcal{P}^{2}_{\succ 0}(^{d_{u}\times d_{y}})$ of probability measures $\pi$ with finite second moment satisfying $\mathcal{C}^{yy}(\pi)\succ 0$.
Throughout, we will denote operators acting on the space of probability measures via the $\mathsf{mathsf}$ font. We introduce the Gaussian projection operator $\mathsf{G}\colon\mathcal{P}(^{n})\to\mathcal{G}(^{n})$ defined by $\mathsf{G}\mu=\mathcal{N}\bigl{(}\mathcal{M}(\mu),\mathcal{C}(\mu)\bigr{)}.$ We refer to $\mathsf{G}$ as a projection because $\mathsf{G}\mu$ is the Gaussian distribution closest to $\mu$ with respect to $\mathrm{KL}(\mu\|\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ [MR2247587], where $\mathrm{KL}(\mu\|\nu)$ denotes the Kullback–Leibler (KL) divergence of $\mu$ from $\nu$. We note that $\mathsf{G}$ defines a nonlinear mapping. Throughout the paper we make use of the following weighted total variation distance as in [carrillo2022ensemble]:
@@
Let $g\colon^{n}\to[1,\infty)$ define the function $g(v):=1+\lvert v\rvert^{2}$. The weighted total variation metric $d_{g}\colon\mathcal{P}(^{n})\times\mathcal{P}(^{n})\to$ is defined by
$d_{g}(\mu_{1},\mu_{2})=\sup_{|f|\leqslant g}\bigl{|}\mu_{1}[f]-\mu_{2}[f]\bigr{|},$
where the supremum is taken over all functions $f\colon^{n}\to$ bounded from above by $g$ pointwise and in absolute value.
@@
  • Unlike the usual total variation distance, the weighted metric in Definition 1.1 enables control of the differences $\lvert\mathcal{M}(\mu_{1})-\mathcal{M}(\mu_{2})\rvert$ and $\lVert\mathcal{C}(\mu_{1})-\mathcal{C}(\mu_{2})\rVert$. This is the content of Lemma B.6, stated in Subsection B.1, which is used in the proof of a key auxiliary result (Lemma B.11).
@@ 1.3 Filtering Distribution
Here we introduce the hidden Markov model that gives rise to the filtering distribution. We employ notation similar to that established in [calvello2025ensemble, carrillo2022ensemble]. We consider $\{u_{j}\}_{j\in\llbracket 0,J\rrbracket}\subset^{d_{u}}$ to be unknown states to be determined from associated observations $\{y_{j}\}_{j\in\llbracket 1,J\rrbracket}\subset^{d_{y}}$. We assume the states and observations to be governed by the following stochastic dynamics and data model:
@@
We assume that the initial state is a Gaussian random variable $u_{0}\sim\mathcal{N}(m_{0},C_{0})$ and that the following independence condition is satisfied:
@@
We define the conditional distribution of the state $u_{j}$ given a realization $Y_{j}^{\dagger}:=\{y_{1}^{\dagger},\dotsc,y_{j}^{\dagger}\}$ as the filtering distribution, which we denote by $\mu_{j}$. We also refer to $\mu_{j}$ as the true filter. In the context of this paper, data $Y_{j}^{\dagger}$ is assumed to be a collection of realizations from (1.4). As formulated in [MR3363508, reich2015probabilistic], the evolution of the filtering distribution may be written as
$\mu_{j+1}=\mathsf{L}_{j}\mathsf{P}\mu_{j},$ (1.6)
where $\mathsf{P}$ and $\mathsf{L}_{j}$ are maps from $\mathcal{P}(^{d_{u}})$ into itself and which effect what are referred to in the data assimilation community as the prediction and analysis steps, respectively [asch2016data]. The prediction operator $\mathsf{P}$ is linear and is determined by the Markov kernel arising from (1.4a). On the other hand, the operator $\mathsf{L}_{j}$ is nonlinear and encodes the incorporation of the new data point $y^{\dagger}_{j+1}$ using Bayes’ theorem. These operators are defined via action on probability measures with Lebesgue density $\mu$ as
$\mathsf{P}\mu(u)$ $=\frac{1}{\sqrt{(2\pi)^{d_{u}}\det\Sigma}}\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\Sigma}^{2}\right)\mu(v)\,\mathrm{d}v,$ (1.7a)
$\mathsf{L}_{j}\mu(u)$ $=\frac{\displaystyle\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u)}{\displaystyle\int_{{}^{d_{u}}}\exp\left(-\frac{1}{2}\bigl{\lvert}y_{j+1}^{\dagger}-h(U)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(U)\,\mathrm{d}U}$ (1.7b)
We may rewrite the analysis map $\mathsf{L}_{j}$ as the composition $\mathsf{B}_{j}\mathsf{Q}$, where the operators $\mathsf{Q}\colon\mathcal{P}(^{d_{u}})\to\mathcal{P}(^{d_{u}}\times^{d_{y}})$ and $\mathsf{B}_{j}\colon\mathcal{P}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ are defined by
$\mathsf{Q}\mu(u,y)$ $=\frac{1}{\sqrt{(2\pi)^{d_{y}}\det\Gamma}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right)\mu(u),$ (1.8a)
$\mathsf{B}_{j}\mu(u)$ $=\frac{\mu(u,y_{j+1}^{\dagger})}{\displaystyle\int_{{}^{d_{u}}}\mu(U,y_{j+1}^{\dagger})\,\mathrm{d}U}.$ (1.8b)
Linear operator $\mathsf{Q}$ maps a probability measure with density $\mu$ into the law of random variable $\bigl{(}U,h(U)+\eta\bigr{)}$, with $U\sim\mu$ independent of $\eta\sim\mathcal{N}(0,\Gamma)$. Nonlinear operator $\mathsf{B}_{j}$ effects conditioning on the datum $y^{\dagger}_{j+1}$. Hence, we may reformulate (1.6) as
@@ 2 The Ensemble Kalman Filter
We begin in Subsection 2.1 by introducing the specific form of the mean field ensemble Kalman filter that we consider in this paper; other versions leading to implementable algorithms and for which similar analysis may be developed can be found in [calvello2025ensemble]. In this subsection we also outline the various assumptions that will be employed for the results of the paper. We note that the subsections that follow proceed with increasingly restrictive assumptions on dynamics and observation operator. However, all of the theorems allow for linear growth of the vector fields defining the dynamics and the observation processes. In Subsection 2.2 we state and prove a stability theorem, which shows that the error between the true filter and the mean field ensemble Kalman filter may be controlled by the error between the true filter and its Gaussian projection on the joint space of state and observations. In Subsection 2.3 we establish an error estimate which shows that the mean field ensemble Kalman filter provides an accurate approximation of the true filtering distribution for dynamics and observation operators that are close-to-linear. In Subsection 2.4 we deploy the results of the two preceding subsections to state a theorem quantifying the error between the finite particle ensemble Kalman filter itself and the true filtering distribution.